# FixAgent: Hierarchical Multi-Agent Framework for Unified Software Debugging

### v2

@@ 4.1 Experimental Setup
Benchmarks. We evaluate FixAgent on four benchmarks featuring bug-fix pairs, including competition programs (Codeflaws Tan et al., 2017, QuixBugs Lin et al., 2017, ConDefects Wu et al., 2024) and real-world projects (Defects4J Just et al., 2014). Codeflaws contains 3902 faulty C programs. QuixBugs includes 40 faulty programs available in both Java and Python. ConDefects recently collected to address data leakage concerns in LLMs, consists of 1,254 Java and 1,625 Python faulty programs. Defects4J, a widely used benchmark from 15 real-world Java projects, features bugs across two versions: version 1.2 with 391 active bugs and version 2.0 adding an additional 415 active bugs, totaling 806. We will report the total number of fixes across these versions.
Baselines. We compare FixAgent against 14 APR baselines, including:
  • Traditional: Angelix (Mechtaev et al., 2016), Prophet (Long, 2018), SPR (Long and Rinard, 2015), and CVC4 (Reynolds et al., 2015).
  • Genetic programming-based: Semfix (Nguyen et al., 2013) and GenProg (Goues et al., 2012; Weimer et al., 2009a).
  • NMT-based: CoCoNuT (Lutellier et al., 2020), CURE (Jiang et al., 2021), and RewardRepair (Ye et al., 2022).
  • Domain knowledge-driven: Tbar (Liu et al., 2019) and Recoder (Zhu et al., 2021).
  • LLM-based (SoTA): AlphaRepair (Xia and Zhang, 2022), Repilot (Wei et al., 2023), and ChatRepair (Xia and Zhang, 2024).
@@
Note that different APR baselines adopted diverse prior fault locations. We use realistic to represent traditional FL and perfect to denote ground-truth FL. We report the results provided in their original papers and follow-up survey Le et al. (2018); Ye et al. (2021); Xia et al. (2023), following previous studies Xia et al. (2023); Xia and Zhang (2024); Lutellier et al. (2020).
@@
LLM Backbones. We apply FixAgent to seven LLMs, including four general-purpose models (gemini-1.5-flash Anil et al., 2023, gpt-3.5-turbo-ca OpenAI, 2023a, gpt-4o OpenAI, 2023b, and claude-3.5-sonnet Anthropic, 2024) and three open-source code LLMs (DeepSeekCoder Guo et al., 2024, CodeLlama-34b Rozière et al., 2023, LLaMA2-70b Touvron et al., 2023). Naturally, these LLMs also serve as baselines for comparison, where we apply the original chain-of-thought (CoT) prompting method Wei et al. (2022). The default backbone of FixAgent is gpt-4o.
@@
Metrics. We utilize the APR metrics, specifically focusing on the number of bugs that are plausibly or correctly fixed. Given the extensive size of ConDefects, we randomize samples from plausible fixes to verify their correctness. Thus, we also employ the metric known as correctness rate—defined as the ratio of correct fixes to plausibly fixed bugs—to assess the effectiveness of FixAgent.
@@
Implementation. For single-file competition programs, only Locator, Fixer, and FixerPro on L2 and L1 need to be initialized, named as FixAgent-Lite. The maximum number of attempts is set to 5 per bug for Lite and 20 for the full version. Previous studies, however, typically sample hundreds to thousands of times. For instance, CoCoNuT samples up to 50,000 patches per bug Lutellier et al. (2020), while ChatRepair samples 100-200 times Xia and Zhang (2024). Our approach is significantly more cost-effective than baselines. Furthermore, we alter the default web search engine (Tavily Tavily, ) to a local one during evaluation to prevent retrieving ground-truth solutions online. The local database consists of the training dataset of CoCoNut (Lutellier et al., 2020) with JavaScript programs removed. Since CoCoNuT was also evaluated on Defects4J and ConDefects, the risk of data leakage is minimal. The static and dynamic analysis is supported by our written scripts and open-source plugins (SonarQube, and GZoltar, ).
@@ 4.2 Comparison with baselines
This section evaluates the debugging capabilities of our framework, FixAgent. The results are shown in Table 1. We do not use ConDefects herein because it is a recently released dataset, and few approaches have been evaluated on it.
Competition Programs. FixAgent plausibly fixes 3130 out of 3982 bugs on Codeflaws, producing 2.2x plausible fixes as the best APR method, GenProg. It has a correctness rate of 95%, i.e., FixAgent correctly repairs 95 bugs out of 100 sampled plausible patches, improving the correctness rate by 60.81% compared to that of CoCoNuT, which correctly fixes the most bugs among APR approaches. FixAgent surpasses the best-performing LLM baseline, claude-3.5-sonnet, by $\sim$19.28with the same correctness rate. Moreover, FixAgent successfully fixes all bugs in QuixBugs across two programming languages, achieving the same SoTA as ChatRepair.
Real-World Software. On Defects4J, FixAgent correctly fixes 197 bugs while plausibly solving 286 bugs, outperforming the SoTA, ChatRepair, by $\sim$25.48%. Moreover, FixAgent correctly fixes 42 unique bugs that the top four baselines have never fixed. Figure 6 shows the Venn diagram of the bugs fixed by the top four baselines and FixAgent on Defects4J. We see that FixAgent correctly fixes 42 unique bugs that these strong baselines have not addressed. We show an example of the unique fixes in A.1.
@@
Overall, LLM-involved baselines perform reasonably well. On top, FixAgent significantly enhances base LLMs on all benchmarks, especially on Defects4J. This is because direct prompting LLMs struggles with too long contexts and complex reasoning. We split debugging into several cognitive steps and introduce external tools and knowledge, thus boosting the capabilities of LLMs.
Takeaway: FixAgent fixes 197 bugs on Defects4J, a 25.48% improvement over the leading baseline. Its lite version fixes all bugs on QuixBugs and achieves 19.28% more plausible fixes on ConDefects with the highest correctness rate.
@@ 4.3 Performance on Different LLMs
To verify the robustness of FixAgent across various LLMs, we compare FixAgent-Lite with its seven LLM backbones, evaluated on 600 randomly sampled bugs from ConDefects (300 for each programming language). Since manually checking the patches is time-consuming, we only present the number of plausible fixes. Table 3 displays the performance gains of FixAgent-Lite (UD-L) over its backbone with CoT prompting. For simplicity, we only report the number of plausible fixes.
Table 3: Performance gains of FixAgent-Lite over different LLMs on 600 samples from ConDefects.
LLMs ConDefects-Java ConDefects-Python CoT UD-L Gain $\uparrow$ CoT UD-L Gain $\uparrow$ CodeLlama-34b 87 113 29.89% 69 86 24.64% LLaMA2-70b 108 147 36.11% 91 133 46.15% DeepSeekCoder 130 198 52.31% 125 178 42.40% gemini-1.5-flash 62 89 43.55% 63 82 30.16% gpt-3.5-turbo-ca 155 191 23.23% 127 174 37.01% claude-3.5-sonnet 213 259 21.60% 186 227 22.04% gpt-4o 211 262 24.17% 179 225 25.70%
The results indicate that FixAgent can consistently enhance its backbone LLM by 21.60%-52.31%. Notably, UD-L with LLaMA2 achieves 280 plausible fixes, closely rivaling the performance of gpt-3.5-turbo-ca (282). Plus, UD-L with DeepSeekCoder plausibly fixes similar bugs (376) as gpt-4o does (390). The results indicate that FixAgent brings the gap between open-source code LLMs and proprietary systems like gpt-4o in debugging. Furthermore, though FixAgent can make improvements across varying LLMs, its overall performance is strongly related to the coding ability of its backbone LLM.
@@ 4.4.1 Impact of Different Agents
To understand the impact of different agents on the effectiveness of FixAgent, we exclude certain agents across Defects4J, as we trigger all agents on it. We also report the number of plausible fixes herein. As indicated by Table 2, the addition of agents different from just Fixer consistently improves the number of plausible fixes. While more agents slightly increase the expenses, the overall performance improves noticeably, demonstrating the effectiveness of the various agents and the importance of the divide-and-conquer idea. Since the higher level is only triggered when the lower level fails, higher levels naturally improve performance.
@@ 5 Conclusion
This paper presents FixAgent, the first end-to-end framework leveraging LLM-based multi-agent synergy to tackle unified software debugging. Our method employs a novel hierarchical coordination paradigm, inspired by a cognitive debugging model, to efficiently manage cognitive steps with minimal communication and dynamically adjust to bug complexity through its three-level architecture. Extensive experiments on four benchmarks demonstrate the superiority of our method over SoTA repair approaches and base LLMs. FixAgent fixes 1.25–2.56$\times$ bugs on a repo-level benchmark and fixes all bugs on QuixBugs. Its lite version achieves the most plausible fixes on the other two competition program benchmarks. Lastly, the effective implementation of Hale and Haworth’s cognitive model for debugging paves new pathways for research in advancing LLM-based multi-agent frameworks for tackling complex coding tasks.
@@ 6 Limitation
While our study introduces the end-to-end framework FixAgent with a hierarchical multi-agent approach for debugging, it does come with certain limitations. First, the complexity of coordinating multiple agents increases computation overhead and can slow down the debugging process, particularly in resource-constrained environments. Second, our framework relies heavily on predefined agent roles and assumptions, which may not fully generalize to novel or edge-case bugs in diverse programming languages. Third, the integration of external program analysis tools, while useful, can introduce delays and inconsistencies in data retrieval, affecting the overall efficiency. Future work should explore optimizing token consumption, improving adaptability to diverse bug types, and ensuring smoother integration with external tools for faster and more reliable debugging.
@@ 7 Ethics Consideration
We do not foresee any immediate ethical or societal risks arising from our work. However, given that FixAgent relies heavily on LLM-generated code patches, there is a potential risk of introducing unintended vulnerabilities or errors in software. We encourage researchers and practitioners to apply FixAgent cautiously, particularly when using it in production environments. Ensuring thorough validation and testing of LLM-generated patches is crucial to mitigate any negative consequences. Additionally, We adhere to the License Agreement of the LLM models and mentioned open-sourced tools.
@@ A.1 Unique Fix Example
We illustrate the power of FixAgent by showing an example bug that is only fixed by FixAgent in Figure 7.
Figure 7: Unique bug fixed by FixAgent in Defects4J.
On the one hand, the fix requires filling in the missing code statements. Traditional FL based on failing test coverage cannot directly identify such an error of lacking the necessary processing of a branching condition. Plus, many template-based or NMT-based APR tools are good at repairing common errors such as syntax errors or simple logic errors, but not at generating new business logic. On the other hand, this fix requires a deep understanding of the specific Jackson deserialization features defined in the package “databind.DeserializationFeature”. Previous LLM-based APR often relies on local context in the prompt and statistical correlation so as to lack the ability to comprehend other code files within the project repo. The three levels of repair enable FixAgent to access extra related information (including templates and repo-level documents) and testing feedback, as well as enhance its reasoning ability under the idea of divide-and-conquer. Combining all these conditions together, FixAgent is able to correctly fix this complex bug.
A.2 Alogrithm of FixAgent
@@
Input: $k$: number of maximum debugging attempts; $m$: number of maximum re-sampling attempts of a single agent; $bug\_meta$: metadata of the bug, a directory including code from the bug-located file, failing test case(s), errors, and program requirements.
@@ A.4 A Demo of the Execution
In this section, we present an example of how FixAgent solves a real-world bug on level-3 repair.
@@ A.4.1 Bug Metadata
Upon receiving the necessary information about a bug, FixAgent initializes agents for problem-solving. Here is an example bug named Lang-1, a real-world bug in Defects4J.
@@
Then, we can roughly locate the buggy code file “NumberUtils.java”, whose code contents are show in Figure 16, as well as the failing oracles in testing code 17.
@@ A.4.2 L3 Repair
For simplicity, we only show the responses from agents on level three of repair. Helper generates a short query summarizing the problem of this bug based on the testing information and buggy code, as shown in Figure 18. All the other agents can benefit from its generated debugging guide. Afterward, RepoFocus lists a list of bug-related files (19). Besides the bug-located file, it also identifies two other files. However, they would not influence the behavior of the number-creation logic unless you are encountering specific exception handling or Unicode string issues when parsing numeric strings. Subsequently, Summarizer generates a code summary for each identified file 20. Thus, we got all the information provided by upstream agents.
@@
Slicer extracts 168 suspicious code lines from 1427 lines in the original buggy code, largely narrowing down the examination scope, as shown in Figure 21. Locator successfully pinpoints the root causes of this bug from the code lines sliced out 22. The following agents can focus on the single logic conditions.

### v3

@@ 4.1 Experimental Setup
Benchmarks. We evaluate UniDebugger on four benchmarks featuring bug-fix pairs, including competition programs (Codeflaws Codeflaws, QuixBugs QuixBugs, ConDefects ConDefects) and real-world projects (Defects4J Defects4J). Codeflaws contains 3902 faulty C programs. QuixBugs includes 40 faulty programs available in both Java and Python. ConDefects recently collected to address data leakage concerns in LLMs, consists of 1,254 Java and 1,625 Python faulty programs. Defects4J, a widely used benchmark from 15 real-world Java projects, features bugs across two versions: version 1.2 with 391 active bugs and version 2.0 adding an additional 415 active bugs, totaling 806. We will report the total number of fixes across these versions.
Baselines. We compare UniDebugger against 14 APR baselines, including:
  • Traditional: Angelix (Angelix), Prophet (Prophet), SPR (SPR), and CVC4 (CVC4).
  • Genetic programming-based: Semfix (SemFix) and GenProg (genProg1; genProg2).
  • NMT-based: CoCoNuT (CoCoNuT), CURE (CURE), and RewardRepair (RewardRepair).
  • Domain knowledge-driven: Tbar (Tbar) and Recoder (Recoder).
  • LLM-based (SoTA): AlphaRepair (AlphaRepair), Repilot (Repilot), and ChatRepair (ChatRepair).
@@
Note that different APR baselines adopted diverse prior fault locations. We use realistic to represent traditional FL and perfect to denote ground-truth FL. We report the results provided in their original papers and follow-up survey Overfitting; QuixBugsSurvey; LLMAPREra, following previous studies LLMAPREra; ChatRepair; CoCoNuT.
@@
LLM Backbones. We apply UniDebugger to seven LLMs, including four general-purpose models (gemini-1.5-flash Gemini, gpt-3.5-turbo-ca GPT3-5, gpt-4o GPT4, and claude-3.5-sonnet Claude) and three open-source code LLMs (DeepSeekCoderV2 DeepSeekCoder, CodeLlama-34b CodeLlama, LLaMA2-70b LLaMA2). Naturally, these LLMs also serve as baselines for comparison, where we apply the original chain-of-thought (CoT) prompting method CoT. The default backbone of UniDebugger is gpt-4o.
@@
Metrics. We utilize the APR metrics, specifically focusing on the number of bugs that are plausibly or correctly fixed. Given the extensive size of ConDefects, we randomize samples from plausible fixes to verify their correctness. Thus, we also employ the metric known as correctness rate—defined as the ratio of correct fixes to plausibly fixed bugs—to assess the effectiveness of UniDebugger.
@@
Implementation. For single-file competition programs, only agents on L2 and L1 need to be initialized, named as UniDebugger-Lite. The maximum number of attempts is set to 5 per bug for Lite and 20 for Full. Each query has a timeout limit of 1 sec. Previous studies typically sample hundreds to thousands of times. For instance, CoCoNuT samples up to 50,000 patches per bug CoCoNuT, while ChatRepair samples 100-200 times ChatRepair. Our approach is significantly more cost-effective than baselines. Furthermore, we alter the default web search engine (Tavily Tavily) to a local one during evaluation to prevent retrieving ground-truth solutions online. The local database consists of the training dataset of CoCoNut (CoCoNuT) with JavaScript programs removed. Since CoCoNuT was also evaluated on Defects4J and ConDefects, the risk of data leakage is minimal. The static and dynamic analysis is supported by our written scripts and open-source plugins (SonarQube and GZoltar).
@@ 4.2 Comparison with baselines
This section evaluates the debugging capabilities of our framework, UniDebugger. The results are shown in Table 1. We do not use ConDefects herein because it is a recently released dataset, and few approaches have been evaluated on it.
Competition Programs. UniDebugger plausibly fixes 3130 out of 3982 bugs on Codeflaws, producing 2.2x plausible fixes as the best APR method, GenProg. It has a correctness rate of 95%, i.e., UniDebugger correctly repairs 95 bugs out of 100 sampled plausible patches, improving the correctness rate by 60.81% compared to that of CoCoNuT, which correctly fixes the most bugs among APR approaches. UniDebugger surpasses the best-performing LLM baseline, claude-3.5-sonnet, by $\sim$19.28% with the same correctness rate. Moreover, UniDebugger successfully fixes all bugs in QuixBugs across two programming languages, achieving the same SoTA as ChatRepair.
Real-World Software. On Defects4J, UniDebugger correctly fixes 197 bugs while plausibly solving 286 bugs, outperforming the SoTA, ChatRepair, by $\sim$25.48%. Moreover, UniDebugger correctly fixes 42 unique bugs that the top four baselines have never fixed. Figure 5 shows the Venn diagram of the bugs fixed by the top four baselines and UniDebugger on Defects4J. We see that UniDebugger correctly fixes 42 unique bugs that these strong baselines have not addressed. We show an example of the unique fixes in A.1.
@@
Overall, LLM-involved baselines perform reasonably well. On top, UniDebugger significantly enhances base LLMs on all benchmarks, especially on Defects4J. This is because direct prompting LLMs struggles with too long contexts and complex reasoning. We split debugging into several cognitive steps and introduce external tools and knowledge, thus boosting the capabilities of LLMs.
Takeaway: UniDebugger fixes 197 bugs on Defects4J, a 25.48% improvement over the leading baseline. Its lite version fixes all bugs on QuixBugs and achieves 19.28% more plausible fixes on ConDefects with the highest correctness rate.
@@ 4.3 Performance on Different LLMs
To verify the robustness of UniDebugger across various LLMs, we compare UniDebugger-Lite with its seven LLM backbones, evaluated on 600 randomly sampled bugs from ConDefects (300 for each programming language). Since manually checking the patches is time-consuming, we only present the number of plausible fixes. Table 3 displays the performance gains of UniDebugger-Lite (UD-L) over its backbone with CoT prompting. For simplicity, we only report the number of plausible fixes.
Table 3: Performance gains of UniDebugger-Lite over different LLMs on 600 samples from ConDefects.
LLMs ConDefects-Java ConDefects-Python
CoT UD-L Gain $\uparrow$ CoT UD-L Gain $\uparrow$
CodeLlama-34b 87 113 29.89% 69 86 24.64%
LLaMA2-70b 108 147 36.11% 91 133 46.15%
DeepSeekCoderV2 130 198 52.31% 125 178 42.40%
gemini-1.5-flash 62 89 43.55% 63 82 30.16%
gpt-3.5-turbo-ca 155 191 23.23% 127 174 37.01%
claude-3.5-sonnet 213 259 21.60% 186 227 22.04%
gpt-4o 211 262 24.17% 179 225 25.70%
The results indicate that UniDebugger can consistently enhance its backbone LLM by 21.60%-52.31%. Notably, UD-L with LLaMA2 achieves 280 plausible fixes, closely rivaling the performance of gpt-3.5-turbo-ca (282). Plus, UD-L with DeepSeekCoderV2 plausibly fixes similar bugs (376) as gpt-4o does (390). The results indicate that UniDebugger brings the gap between open-source code LLMs and proprietary systems like gpt-4o in debugging. Furthermore, though UniDebugger can make improvements across varying LLMs, its overall performance is strongly related to the coding ability of its backbone LLM.
@@ 4.4.1 Impact of Different Agents
To understand the impact of different agents on the effectiveness of UniDebugger, we exclude certain agents across Defects4J, as we trigger all agents on it. We also report the number of plausible fixes herein. As indicated by Table 2, the addition of agents different from just Fixer consistently improves the number of plausible fixes. While more agents slightly increase the expenses, the overall performance improves noticeably, demonstrating the effectiveness of the various agents and the importance of the divide-and-conquer idea. Since the higher level is only triggered when the lower level fails, higher levels naturally improve performance.
@@ 5 Conclusion
[…]
@@ 6 Limitation
Traditional APR tasks typically assume the presence of failing test cases to guide fault localization and repair, which aligns with the CI/CD pipeline where developers encounter bugs during automated testing. In such scenarios, the debugging process relies on concrete evidence (e.g., test failures, stack traces) to isolate errors. UniDebugger is specifically designed for this context, where test cases serve as the primary oracle to validate fixes. On the one hand, this design mirrors real-world developer workflows, where unit tests and integration tests are integral to identifying and resolving bugs during development. On the other hand, we follow previous studies APRQuality; Recoder; Tbar; ChatRepair of modeling the APR task in a test-driven framework.
In contrast, issue-driven repair (e.g., SWE-bench SWE-Bench) targets user-reported bugs described in natural language, often lacking explicit test cases or formal specifications. While this scenario is practically relevant, it emphasizes replicating and fixing bugs based on user observations (e.g., "the application crashes when clicking button X"), while test-driven APR focuses on resolving errors detectable for developers before software release. The latter provides a structured, reproducible environment for evaluating automated debugging systems. However, extending UniDebugger to issue-driven contexts would require integrating natural language understanding modules and replicating user-described failures, which remains a direction for future work. Furthermore, although this debugging model falls within a well-studied scope of program repair, the ability of our approach to address broader classes of bugs, such as configuration errors, remains unknown. Future work should explore optimizing token consumption, improving adaptability to diverse bug types, and ensuring smoother integration with external tools for faster and more reliable debugging.
@@ 7 Ethics Consideration
We do not foresee any immediate ethical or societal risks arising from our work. However, given that UniDebugger relies heavily on LLM-generated code patches, there is a potential risk of introducing unintended vulnerabilities or errors in software. We encourage researchers and practitioners to apply UniDebugger cautiously, particularly when using it in production environments. Ensuring thorough validation and testing of LLM-generated patches is crucial to mitigate any negative consequences. Additionally, We adhere to the License Agreement of the LLM models and mentioned open-sourced tools.
8 Artifact Discussion
All adopted benchmarks, LLM models, and open-source tools are used strictly within their intended research purposes as defined by their creators:
  • Benchmark datasets (Codeflaws, QuixBugs, ConDefects, Defects4J) are all academic resources explicitly created for evaluating program repair techniques.
  • Baseline APR tools (e.g., Angelix, Semfix, CoCoNuT) are used in accordance with their original research implementations
  • LLM backbones (gpt-4o, claude-3.5-sonnet, etc.) are accessed through official APIs under research-only agreements.
  • Code analysis tools (SonarQube, GZoltar) are used as open-source static analysis utilities per their LGPL licenses.
Acknowledgments
This work was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14209124 of the General Research Fund).
@@ A.1 Unique Fix Example
We illustrate the power of UniDebugger by showing an example bug that is only fixed by UniDebugger in Figure 6.
Figure 6: Unique bug fixed by UniDebugger in Defects4J.
On the one hand, the fix requires filling in the missing code statements. Traditional FL based on failing test coverage cannot directly identify such an error of lacking the necessary processing of a branching condition. Plus, many template-based or NMT-based APR tools are good at repairing common errors such as syntax errors or simple logic errors, but not at generating new business logic. On the other hand, this fix requires a deep understanding of the specific Jackson deserialization features defined in the package “databind.DeserializationFeature”. Previous LLM-based APR often relies on local context in the prompt and statistical correlation so as to lack the ability to comprehend other code files within the project repo. The three levels of repair enable UniDebugger to access extra related information (including templates and repo-level documents) and testing feedback, as well as enhance its reasoning ability under the idea of divide-and-conquer. Combining all these conditions together, UniDebugger is able to correctly fix this complex bug.
A.2 Alogrithm of UniDebugger
@@
Input: $k$: number of maximum debugging attempts; $m$: number of maximum re-sampling attempts of a single agent; bug_metabug\_meta: metadata of the bug, a directory including code from the bug-located file, failing test case(s), errors, and program requirements.
@@ A.4 A Demo of the Execution
In this section, we present an example of how UniDebugger solves a real-world bug on level-3 repair.
@@ A.4.1 Bug Metadata
Upon receiving the necessary information about a bug, UniDebugger initializes agents for problem-solving. Here is an example bug named Lang-1, a real-world bug in Defects4J.
@@
Then, we can roughly locate the buggy code file “NumberUtils.java”, whose code contents are show in Figure 15, as well as the failing oracles in testing code 16.
@@ A.4.2 L3 Repair
For simplicity, we only show the responses from agents on level three of repair. Helper generates a short query summarizing the problem of this bug based on the testing information and buggy code, as shown in Figure 17. All the other agents can benefit from its generated debugging guide. Afterward, RepoFocus lists a list of bug-related files (18). Besides the bug-located file, it also identifies two other files. However, they would not influence the behavior of the number-creation logic unless you are encountering specific exception handling or Unicode string issues when parsing numeric strings. Subsequently, Summarizer generates a code summary for each identified file 19. Thus, we got all the information provided by upstream agents.
@@
Slicer extracts 168 suspicious code lines from 1427 lines in the original buggy code, largely narrowing down the examination scope, as shown in Figure 20. Locator successfully pinpoints the root causes of this bug from the code lines sliced out 21. The following agents can focus on the single logic conditions.